OpenStudy (anonymous):

Which expressions, when evaluated with y = –7, are positive in value? Choose all answers that are correct. A. 18 + 2y B. –3y + (–17) C. –2y – 23 D. –30 – 4y

OpenStudy (anonymous):

@Conqueror @confluxepic

sammixboo (sammixboo):

Here let's plug in -7 where it belongs in each of the equation in your choices :) A. \(18~ +~ 2\times (-7)\) B. \(-3\times (-7)~+~ (–17)\) C. \(-2\times (-7)~-~23\) D. \(-30 ~- ~4~\times~(-7)\) Let's solve Option A first :) \(18~ +~ 2\times (-7)\) First we must multiply, and the only multiplication in this problem is \( 2\times (-7)\), so what is \(2\times (-7)\)?

sammixboo (sammixboo):

(refresh if you see any black diamonds with question marks) :)

sammixboo (sammixboo):

Are you there?